OpenStudy (anonymous):

*Helper will recieve a medal :)* identify conesection and state important parts y^2=x+16y

OpenStudy (accessdenied):

What are considered "important parts"? Is that: vertex, focus, directrix any others we need to concern ourselves with?

OpenStudy (anonymous):

vertices,co-vertices, foci, radius center depends on what it is?

OpenStudy (accessdenied):

Oh got it. I had already identified this as a parabola because only the y was squared. There is no squared term for x, so it does not become an ellipse/circle or hyperbola.

OpenStudy (accessdenied):

Parabolas always have the form: y = ax^2 or x = ay^2 barring any extras such as vertex shifts

OpenStudy (anonymous):

alright

OpenStudy (accessdenied):

So we should solve this equation for x first: y^2 = x+16y and then try to put it into vertex form of a parabola by completing the square with y^2 - 16y. You know how to do that part?

OpenStudy (anonymous):

would it be y^2-16y+64?

OpenStudy (accessdenied):

Yup. y^ 2 - 16y + 64 - 64 = x

OpenStudy (anonymous):

then that would be (y-8)^2=x

OpenStudy (accessdenied):

There will still be the - 64 part because otherwise we have changed the equation' s value (y^ 2 - 16y + 64) - 64 = x (y - 8)^2 - 64 = x

OpenStudy (anonymous):

now how do you find the vertex, focus and directrix?

OpenStudy (accessdenied):

Vertex is the easiest to find first. Starting with the equation with the vertex at (0, 0) and applying some transformations to get it to match our current equation... y^2 = x We shifted the y-value UP 8 units, which makes for subtracting (it takes greater y-values to reach the same point now) (y - 8)^2 = x and then shifted it to the LEFT 64 units: (y - 8)^2 - 64 = x So our vertex has moved: (0 - 64, 0 + 8) = (-64, 8)

OpenStudy (accessdenied):

The focus and directrix of the parabola are subject to the same translations. x = ay^2 The distance from vertex to directrix or focus is c= 1/4a The orientation of the parabola puts the focus in front of (horizontally) the vertex, and the directrix vertical behind the parabola: